Find the measure of Angle 3.

measure of angle 1 = x
measure of angle 2 = x-6

2_13c
1
Expert's answer
2019-06-10T13:08:23-0400

Each triangle has an angle sum of 180 degrees.

Thus, α1 + α2 + α3 = 180°

α3 = 180° - α2 - α1 = 180° - x - (x-6°) = 186° - 2x